# An Exposition of Pathfinding Strategies Within Lightning Network Clients

### v1

@@ VII Evaluation
LND1’s low success rate can be attributed to the apriori estimation of the success probability, which assumes a base success likelihood of 0.6. This assumption, along with default values for the constants $\mathrm{c_{o}}$ and $\mathrm{s_{o}}$ in (13), might not yield reliable paths with. Moreover, this model takes into account historical payment attempts to adjust the channel-wise success probability as described in Equation 9. However, in our simulation, we did not consider historical attempts when computing the success probabilities, which might have contributed to the low success rate.
Examination of the weight function (32) employed by Eclair_case3 reveals a tendency to favor paths with lower fees, as evidenced by the relatively diminished influence of other design elements within the weight function, which are assigned lower cost, and consequently contribute less weight compared to fees. This trend is consistent with the observed outcomes, indicating that Eclair_case3 consistently offers paths characterized by lower fees.
The efficacy of CLN in providing paths with low timelock can also be attributed to the design of its weight function, which prioritizes paths with lower timelock. Additionally, CLN’s competitive performance with poorly connected nodes can be linked to $\mathrm{capacityBias}_{e}$ in (17), as it penalizes channels where a significant portion of channel capacity is utilized. This design feature enables CLN to perform more effectively among nodes with large capacity. In our experimental setup, some poorly connected nodes possess channels with substantial capacity, CLN’s weight function incentivizes such channels. Consequently, resulting in competitive performance in such scenarios.
Finally, LDK’s decline in reliability for larger amounts can be attributed to the weight function (19) which prioritizes paths with lower fees over higher reliability. The higher fee ratio for small payment amounts in LDK may result from the weight function favoring paths with lower HTLC minimums rather than lower fees as illustrated in Section VI-C.
@@ VIII Limitations
In our simulation model, we do not consider the historical payment attempts to update the channel-wise success probability. In the event of failure, the success probability of the failed channel is not taken to zero, as it is in LND1.
Furthermore, our model neither has the capability to record historical liquidity estimations like channel upper bound and lower bounds, as in LDK, nor does it have the capability to record the previous success and failure amounts in the channel, as is the case in LND.
Our model does not consider all the side constraints during pathfinding. Specifically, we do not consider constraints on timelock, path length, and fees.
This limitation may introduce slight variations in the results observed in real Lightning Network routing. Additionally, we used default values for design elements that are constants. Varying these values could yield different results.
@@ IX Conclusion
In this study, we assess the empirical performance of four prominent Lightning Network clients: LND, CLN, LDK, and Eclair, across various metrics including success rate, fee ratio, path length, and timelock. Our experiments reveal that LND demonstrates superior success rates, which can be attributed to its detailed modeling of success probabilities and its larger emphasis on these in the implied path cost. Conversely, LDK demonstrates moderate reliability but suboptimal fee performance for smaller payments, while Eclair distinguishes itself by offering paths with low fees. Additionally, CLN stands out for providing paths with minimal timelock, which is aligned with its weight function modelling which is designed to prioritize timelock minimization.
We find that the connectivity level of nodes significantly influences the selection of an appropriate path. Depending on the node’s connectivity level and the desired performance metric optimization, the selection of the appropriate weight function or routing client becomes crucial.
Furthermore, we observe that certain path cost modeling, while reasonable from a modeling perspective, can harm the performance of the pathfinding algorithm, and simple, modifications to standard shortest path algorithms such as Dijkstra’s algorithm do not always return the expected, optimal solution, in particular, in the case of LND’s cost function. We also observe that the imposition of side constraints (see (8) and Table I) imposes potentially severe difficulties for the pathfinding process, as the underlying problem becomes a constrained shortest-path problem, which is $\mathcal{NP}$-complete and thus cannot be guaranteed to be solved optimally in practice for all but small graph instances.

### v2

@@ VII-C Analysis of Results
LND-ap’s lower success rate can be attributed to the $Apriori$ estimation of the success probability, which assumes a base success likelihood of 0.6. This assumption, along with default values for the constants $\mathrm{c_{o}}$ and $\mathrm{s_{o}}$ in (16), might not yield reliable paths. Moreover, this model takes into account historical payment attempts to adjust the channel-wise success probability as described in Equation 13. However, in our simulation, we do not consider historical attempts when computing the channel success probabilities, which might have contributed to the low success rate.
By analyzing the different variants of Eclair’s pathfinding module, we see that the weight function (37) used by Eclair3 reveals a tendency to favor paths that balance high reliability and low fees. This trend is consistent with the observed outcomes, indicating that Eclair3 consistently attempts to provide paths characterized by both higher reliability and lower fees. Finally, the decline in reliability for larger amounts in both LDK-un and LDK-bm can be attributed to the weight function (22), which prioritizes paths with lower fees over higher reliability. The higher fee ratio for small payment amounts in LDK may result from the weight function favoring paths with lower HTLC minimums rather than lower fees as illustrated in Section VI-C. Moreover, LDK employs historical liquidity tracking to estimate upper and lower bounds of channel liquidity. Since our simulations do not incorporate historical data, this omission may have contributed to the low success rates observed in LDK. Incorporating historical data has a clear potential to further improve its success rates.
@@ VIII Limitations
In our simulation model, we do not consider the historical payment attempts to update the channel-wise success probability. In the event of failure, the success probability of the failed channel is not taken to zero, as it is in LND-ap.
Furthermore, our model neither has the capability to keep track of historical liquidity estimations like channel upper bound and lower bounds, as in LDK-un and LDK-bm, nor does it have the capability to record the previous success and failure amounts in the channel, as is the case in LND-bm.
Our model does not consider most side constraints (see Table I) during pathfinding, as these change the complexity class of the pathfinding problem, cf. Section V. Specifically, we do not consider constraints on timelock, path length, and fees.
This limitation may introduce slight variations in the results observed in real Lightning Network routing. Additionally, we used default values provided by the LN clients’ open-source implementations for design elements that are constants. Varying these values could yield different results.
@@ IX Conclusion
In this study, we provide transparency into the pathfinding modules of the four prominent Lightning Network clients LND, CLN, LDK, and Eclair. Using simulated channel liquidities and payment attempts, we further assess their empirical performance across various metrics including success rate, fee ratio, path length, and timelock. Our experiments reveal that Eclair demonstrates superior success rates and distinguishes itself by offering paths with low fees. This performance can be attributed to its weight function, which is designed to trade off between high reliability and low fees. Although LND provides moderate reliability with its default settings, it achieves higher success rates when these settings are fine-tuned. LDK exhibits suboptimal fee performance for smaller payments; however, for larger payment amounts, it provides paths with the lowest fees. Additionally, CLN stands out for offering paths with minimal timelock, consistent with its weight function, which is designed to prioritize timelock minimization.
We find that the connectivity level of the nodes significantly influences the selection of an appropriate path. Depending on the node’s connectivity level and the desired performance metric optimization, the selection of the appropriate weight function or routing client becomes crucial.
Furthermore, we observe that certain path cost modeling, while reasonable from a modeling perspective, can harm the performance of the pathfinding algorithm, and simple modifications to standard shortest path algorithms such as Dijkstra’s algorithm do not always return the expected, optimal solution, in particular, in the case of LND’s cost function. We also observe that the imposition of side constraints (see (12) and Table I) imposes potentially severe difficulties for the pathfinding process, as the underlying problem becomes a constrained shortest-path problem, which is $\mathcal{NP}$-complete and thus cannot be guaranteed to be solved optimally in practice for all but small graph instances.
@@
Acknowledgements
The authors thank René Pickhardt, Matt Corallo (BlueMatt) and bitromortac for their helpful comments and feedback on our manuscript. The authors are supported by the National Science Foundation under the grant NSF-CIF-2348430.